OpenStudy (anonymous):

Will give Medal and Fan. f(x)=(x^2+3x-4) and g(x)=(x+4). what is f*g and it's domain. What is f/g and it's domain.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@uri

OpenStudy (campbell_st):

well if you write it out\[\frac{f(x)}{g(x)} = \frac{(x^2 + 3x -4)}{x - 4}\] you might like to factor the numerator.... and see if there are common factors.

OpenStudy (anonymous):

Well the g(x)=x+4

OpenStudy (campbell_st):

opps yes the denominator should be (x + 4)

OpenStudy (anonymous):

So I can remove one x correct?

OpenStudy (campbell_st):

no... factor the numerator..... can you do that...?

OpenStudy (anonymous):

Well, -3 and -1

OpenStudy (anonymous):

I suck at anything to do with factoring, so Helping me with this first part will be the hardest part I promise lol.

OpenStudy (campbell_st):

not quite... find the factors of -4 that add to 3.... the larger factor is positive and the smaller factor is negative.

OpenStudy (anonymous):

so 4 and -1, yeah. I did it the opposit way and found things that add to -4 and multiply to 3.

OpenStudy (campbell_st):

so the numerator is (x + 4)(x -1) is there a common factor...?

OpenStudy (anonymous):

4...?

OpenStudy (campbell_st):

well the common factor is (x + 4) which means f/g = (x - 1) so there are 2 options... the simplified verion has a domain of all real x. of if you look at the original problem x cannont be -4 as this gives a zero denominator. the point x = -4 in this case is a point of discontinuity so it depends on the level of math you are doing.... choices, all real x or all real x except x = -4

OpenStudy (anonymous):

Okay. I understand, after the factoring incident lol. All real x except x=-4 is what i was looking for. Thank you.